OpenStudy (anonymous):

the perimeter of a rectangle is 80x.the base is 7 times its height in terms of x , what are the demension of the rectangle

OpenStudy (anonymous):

Let height be x then base is 7x Now Permiter = 2(height + base) therefore, 80x = 2(7x + x) Solve for x

OpenStudy (anonymous):

@Soumyadeep According to you response, 80x = 16x

OpenStudy (anonymous):

your*

OpenStudy (anonymous):

Let's call the height of the rectangle a. B= 7a, H= a 2(B+H)= 2(7a+a)=16a; 16a=80x. Divide both sides by 16, so that we get a = 5x. B=7a=7(5x)=35x, H=a=1(5x)=5x.